OpenStudy (anonymous):

How many pounds of cashew nuts worth 75 cents a pound must be mixed with 10 pounds of pecans worth $1.50 a pound to make a mixture worth $1.00 a pound? Which of the equations could be used to solve for p, the pounds of cashew nuts? 75p + 150(10) = 100(p + 10) 75(p + 10) + 150(10) = 100(p + 10) 75p + 150 = 100(p + 10)

OpenStudy (anonymous):

.75(m)+10(1.50)=1(m+10) .75m+15=m+10 .25m=5 m=5/.25=20 You need 20 lbs. of cashew nuts..